# Ground-state phase diagram of $S=1/2$ Heisenberg model on 2D square-hexagon-octagon lattice

### v1

@@ Abstract
Using stochastic series expansion quantum Monte Carlo method and density matrix renormalization group, we study the ground-state phase diagram of $S=1/2$ Heisenberg model on 2D square-hexagon-octagon (SHO) lattice. In this model, we consider two kinds of nearest-neighbor interaction (intra-hexagon interaction $J_{1}$ and inter-hexagon $J_{2}$) and the selected third nearest-neighbor interaction $J_{3}$ along $x$ direction. From our calculations, there are five phases in the parameters regime $0<\lambda_{1}=J_{2}/J_{1}<4,0<\lambda_{2}=J_{3}/J_{1}<4$, including a Néel antiferromagentic phase, a Haldane-like symmetry protected topological phase, a hexagon phase and two dimer phases. In the Haldane-like SPT phase, we characterized its topological nature using the degeneracy of ground-state energy under open boundary condition and the entanglement spectrum. To characterize the phase boundaries, we use spin stiffness and Binder cumulant to do the comprehensive finite-size scalings. From data collapse, the critical behaviors of all the nonmagnetic phases to the antiferromagnetic phase belong to the 3D $O(3)$ Heisenberg universality class. As a theoretical exploration, our work establishes a foundational framework for understanding 2D magnetism on the SHO lattice.
@@ I Introduction
Since the mechanical exfoliation of a single layer of graphene in 2004 [1, 2], two-dimensional (2D) material systems have garnered significant attention [3, 4, 5, 6]. These breakthroughs has spurred extensive research into the properties and applications of 2D materials. Graphene is a periodic hexagonal network built from $sp^{2}$ -hybridized carbon atoms, with excellent electrical and mechanical properties. Based on the structure of graphene, researchers try to obtain two-dimensional materials with some new properties by adding non-hexagons to replace six-membered rings in graphene. For example, a new kind of Graphene-like nanoribbons periodically embedded with four- and eight-membered rings has been generated [7]. In this material, the non-hexagonal rings alter the electronic properties of the nanoribbons due to the topological change. Similar to graphene nanoribbons, 2D nonbenzenoid carbon allotropes have also attracted great interest from researchers. In 2021, an article reported a 2D biphenylene network composed of $sp^{2}$-hybridized carbon atoms with periodically arranged four-, six-, and eight-membered rings [8], achieving a two-dimensional square-hexagon-octagon (SHO) lattice, as illustrated in Fig. 1. The SHO lattice as a new kind of unfrustrated 2D lattice offers a new platform for studying electronic correlations, magnetism, and superconducting properties.
@@
Figure 1: (a)The sketch of SHO structure, constituted with hexagon unit cells. There are three types of bonds with different interactions from our definition: $J_{1}$ indicated with black color, $J_{2}$ is red and $J_{3}$ is blue, respectively. We also define the ratios between them: $\lambda_{1}=J_{2}/J_{1}$ and $\lambda_{2}=J_{3}/J_{1}$. A unit cell of SHO lattice consists of six sites, which is shown in the black dashed square. (b) The lattice in square form which is topologically equivalent to the SHO lattice.
@@
The electronic band structures were systematically investigated using density functional theory (DFT) [9, 10, 11], and the possible superconductivity properties of biphenylene were studied using the random phase approximation (RPA) [10]. Furthermore, the ground-state phase diagram and critical behavior of the Hubbard model on the SHO lattice were examined using the determinant quantum Monte Carlo method [12]. Additionally, the topological and electronic properties on an anisotropic SHO lattice were also studied [13]. However, the magnetic properties of the Heisenberg model on the SHO lattice remain an open question. It is of great interest to investigate the ground-state properties of the $S=1/2$ Heisenberg model on the SHO lattice. Although no SHO lattice magnetic materials have been identified in transition metals thus far, there is potential for their discovery in the future. Our research offers crucial theoretical analysis in anticipation of such findings. Moreover, magnetic systems of this type could potentially be realized in experimental platforms such as cold atoms or metal-organic frameworks (MOFs). As a new theoretical model, it is of significant value for the systematic investigation of its rich phase diagram and critical behavior.
@@
In this work, we employ the stochastic series expansion (SSE) quantum Monte Carlo (QMC) [14, 15, 16, 17] and density matrix renormalization group (DMRG) [18, 19, 20, 21, 22, 23] to study the ground-state phase diagram of $S=1/2$ Heisenberg model on the SHO lattice. To enrich the phase diagram and enhance its topological equivalence to the square lattice, we introduce the selected third nearest-neighbor interaction $J_{3}$ along $x$ direction. Our findings include a Néel antiferromagnetic phase, two dimerized phases, a hexagon phase and a Haldane-like phase. The Haldane-like phase [24, 25] has a unique ground state which does not break any symmetries and exhibits a finite triplet excitation gap. Its topological properties can be evidenced by the degeneracy of the ground-state energy levels under open boundary condition and the two-fold degeneracy of the entanglement spectrum. This phase is distinct and cannot be continuously connected to trivial phase in the presence of certain symmetries, making it a symmetry-protected topological phase [26, 27, 28, 29]. Similar phases have been observed in $S=1/2$ two-leg ladder systems [30, 31, 32]. To obtain accurate phase boundaries and critical behaviors between different phases, We also use comprehensive finite-size scaling techniques [33, 34] to verify the universality class.
@@
The paper is organized as follows. In Sec. II, we present the Hamiltonian and the physical quantities used to investigate the ground-state phase diagram of the SHO lattice. In Sec. III, we employ SSE-QMC and DMRG methods with detailed finite-size scaling to numerically determine the accurate ground-state phase diagram. Specifically, Subsec. III.1 examines the case of $\lambda_{2}=J_{3}/J_{1}=0$, Sec. III.2 considers the condition of $\lambda_{1}=J_{2}/J_{1}=0$, and Sec. III.3 explores the full phase diagram. Additionally, Sec. III.4 focuses on the critical behavior of phase transitions in the SHO lattice. Finally, in Sec. IV, we summarize the identified ground - state phases and provide a brief discussion of them.
@@ II MODEL AND METHODS
The Hamiltonian of $S=1/2$ Heisenberg model on SHO lattice can be written as the following form,
@@
$+J_{3}\sum_{\langle\langle\langle i,j\rangle\rangle\rangle}\mathbf{S}_{i}\cdot\mathbf{S}_{j},$ (1)
where $\langle i,j\rangle$ and $\langle i,j\rangle^{\prime}$ denote the intra- and inter-hexagon nearest-neighbor interaction, respectively, $\langle\langle\langle i,j\rangle\rangle\rangle$ denotes the selected third nearest-neighbor interaction along $x$ direction, and $S_{i}$ denotes the spin operator at site $i$. For simplicity, we set the inter-hexagon interactions to be isotropic (i.e., the horizontal and vertical $J_{2}$ bonds are equal). Furthermore, the $J_{3}$ coupling was introduced specifically to establish a connection with the square-lattice Heisenberg model. To compare the relative strength of interactions, we set $\lambda_{1}=J_{2}/J_{1},\lambda_{2}=J_{3}/J_{1}$.
In the following text, unless specified, we generally employ periodic boundary conditions in QMC calculations. The linear system size is denoted as $L$, and the total number of grid points is given by $N=6L^{2}$. In the SSE-QMC calculation, We use the staggered magnetization along $S^{z}$ direction and the squared staggered magnetization $M_{s}^{2}$ to illustrate the Néel order:
@@
$M_{s}^{2}=$ $\frac{1}{N^{2}}\sum_{ij}\phi_{i}\phi_{j}\langle\mathbf{S}_{i}\cdot\mathbf{S}_{j}\rangle$ (3)
@@
$=\frac{3}{2\beta N}\left\langle\left(N_{a}^{+}-N_{a}^{-}\right)^{2}\right\rangle,\quad a=x,y$ (5)
@@
Spin stiffness denotes the ground state energy cost of applying a twisting angle $\theta$ to the rotors of a spin system, which can be calculated by the number of $S_{i}^{+}S_{j}^{-}$ and $S_{i}^{-}S_{j}^{+}$ operators in SSE program. Here $E_{0}$ is the ground state energy of the twisted Hamiltonian, $N$ is the number of spins in the system, $N_{a}^{+}$ is the number of $S_{i}^{+}S_{j}^{-}$ operators, and $N_{a}^{-}$ is the number of $S_{i}^{-}S_{j}^{+}$ operators. The value of $\beta=1/T$ is taken to be proportional to the system size $L$ when doing the scaling. The factor $3/2$ in the Binder cumulant is for normalization, i.e. $U_{2}\rightarrow 1$ when the system being a magnetic ordered state and $U_{2}\rightarrow 0$ when being a disordered state, for $N\rightarrow\infty$.
When $J_{2}=0$, the system becomes a ladder system. To analyze the ground-state phases in this scenario, we use the DMRG method to compute the energy gaps and entanglement spectra. In our DMRG calculations, we use 4000 $SU(2)$ symmetric states to ensure that the density matrix truncation error is below $10^{-8}$. The energy gaps are defined as follows: The singlet gap $\Delta_{S}$ is the energy difference between the first excited state and the ground state in the $S=0$ sector, The triplet gap $\Delta_{T}$ is the energy difference between the lowest $S=1$ state and the $S=0$ ground state. The quintuplet gap $\Delta_{Q}$ is the energy difference between the lowest $S=2$ state and the $S=0$ ground state, i.e.,
@@ III NUMERICAL RESULTS
Using SSE-QMC method and finite-size scaling, we obtain the accurate ground-state phase diagram of the $S=1/2$ Heisenberg model on 2D SHO lattice, which is shown in Fig. 2. There are four gapped phases, namely the hexagon phase, the orthogonal staggered dimer (OSD) phase, the ladder staggered dimer (LSD) phase and the Haldane-like symmetry protected topological (SPT) phase, surrounding the gapless Néel antiferromagnetic (AFM) phase. In the subsequent subsection, we will elaborate on the method used to obtain the phase diagram and provide a detailed overview of the distinct phases.
@@
Figure 2: The ground-state phase diagram of $S=1/2$ Heisenberg model on 2D SHO lattice. In total there are five phases in the phase diagram, namely the hexagon phase, the orthogonal staggered dimer (OSD) phase, the ladder staggered dimer (LSD) phase, the Haldane-like symmetry protected topological (SPT) phase, and the Néel antiferromagnetic (AFM) phase.
III.1 $J_{1}$-$J_{2}$ ground-state phase diagram
@@
First, we investigate the ground state of the system with $\lambda_{2}=J_{3}/J_{1}=0$ fixed and a changing $\lambda_{1}=J_{2}/J_{1}$. Initially, we examined the ground state of the Heisenberg model on square-hexagon-octagon lattice at two extremes: $\lambda_{1}$ approaching 0 and positive infinity. At $\lambda_{1}=\lambda_{2}=0$, the model reduces to isolated, orderly hexagonal lattices. The ground state here is a direct product of hexagonal singlet states with finite local excitation gaps. After adding weak inter-hexagon interactions, the system is in a hexagon phase with gapped lowest triplet excitation. When $\lambda_{1}\to\infty$ and $\lambda_{2}=0$, the ground state adiabatically connects to a vertically and horizontally regular dimer product state, termed the orthogonal staggered dimer (OSD) phase.
@@
Regarding the phase between the hexagon phase and the OSD phase, given the even-number sublattice structure of the SHO lattice, it is possible that a long-range magnetic order exists. And it is tentatively speculated to be the Néel antiferromagnetic (AFM) phase. By extrapolating the squared magnetization at $\lambda_{1}=1$ (as shown in Fig. 3 (b)), we can confirm that the intermediate phase is indeed an Néel AFM phase. To characterize the phase boundaries among the three phases, we show the intersection of Binder cumulant $U_{2}$ for different sizes when $\lambda_{1}\in[0,3]$, which is illustrated in Fig. 3 (a). To pinpoint the crossing location precisely, we computed additional, denser data points in the vicinity of the intersection, enabling a precise determination of the crossing point, as shown in the inset of Fig. 3 (a). Subsequently we perform an extrapolation of this intersection to the thermodynamic limit of $1/L\to 0$ using scaling function $\lambda_{1,c}(L)=\lambda_{1,c}(\infty)+aL^{-\omega}$ [14]. Based on the finite-size scaling results, it is evident that the system undergoes phase transitions at $\lambda_{1,c}=0.664(1)$ and $\lambda_{1,c}=1.379(1)$ in the thermodynamic limit.
@@
Figure 3: (a)Variation of Binder cumulant $U_{2}$ with $\lambda_{1}$ at different sizes when $\lambda_{2}=0$. Inset in (a) is the extrapolation of the two intersection positions of the Binder cumulant as a function of the size, using the extrapolation function in the form of $\lambda_{1,c}(L)=\lambda_{1,c}(\infty)+aL^{-\omega}$ [14]. (b) Second-order polynomial extrapolation of $M_{s}^{2}$ as a function of $L$ when $\lambda_{1}=1$. The nonzero extrapolated value indicates that there is a nonzero staggered magnetic order in the thermodynamic limit: $M^{2}_{s}(\infty)=0.0609(2)$, and $SU(2)$ continuous symmetry breaking occurs at the intermediate phase.
@@
Hence, when $\lambda_{2}=0$ and $\lambda_{1}$ varies (i.e., in the $J_{1}-J_{2}$ model), the ground-state phase diagram consist of three phases: the hexagon phase, AFM phase, and the OSD phase. The hexagon phase and OSD phase are nonmagnetic, featuring nonzero triplet excitation energy gaps. In contrast, the AFM phase is magnetic, characterized by gapless Goldstone mode excitations.
III.2 $J_{1}$-$J_{3}$ ground-state phase diagram

### v2

# Ground-state phase diagram and Haldane-like SPT phase in the $S=1/2$ Heisenberg model on the square-hexagon-octagon lattice

@@ Abstract
Using stochastic series expansion quantum Monte Carlo and density matrix renormalization group methods, we investigate the ground-state phase diagram of the $S=1/2$ Heisenberg model on the two-dimensional square-hexagon-octagon (SHO) lattice. The model incorporates nearest-neighbor interactions $J_{1}$ (intrahexagon interaction) and $J_{2}$ (interhexagon), as well as a selected third-neighbor interaction $J_{3}$ along the $x$ direction. We identify five distinct phases in the parameter regime $0<\lambda_{1}=J_{2}/J_{1}<4,0<\lambda_{2}=J_{3}/J_{1}<4$: a Néel antiferromagentic phase, two dimer phases (orthogonal and ladder staggered dimers), a hexagon singlet phase, and notably a Haldane-like symmetry-protected topological (SPT) phase. The topological nature of the Haldane-like phase is confirmed by the degeneracy of the ground-state energy under open boundary conditions and the twofold degeneracy of the entanglement spectrum. Phase boundaries are accurately determined using finite-size scaling of the spin stiffness and Binder cumulant. Data collapse analysis reveals that all transitions from nonmagnetic phases to the antiferromagnetic phase belong to the three-dimensional $O(3)$ Heisenberg universality class. In addition, we investigate the robustness of the SPT phase to other interactions, such as those that could arise in experimental materials. Our work establishes a comprehensive theoretical framework for understanding magnetic and topological phases on the SHO lattice.
@@ I Introduction
Since the mechanical exfoliation of a single layer of graphene in 2004 [1, 2], two-dimensional (2D) material systems have garnered significant attention [3, 4, 5, 6]. These breakthroughs has spurred extensive research into the properties and applications of 2D materials. Graphene is a periodic hexagonal network built from $sp^{2}$-hybridized carbon atoms, with excellent electrical and mechanical properties. Based on the structure of graphene, researchers try to obtain two-dimensional materials with some new properties by adding nonhexagons to replace six-membered rings in graphene. For example, a new kind of graphenelike nanoribbons periodically embedded with four- and eight-membered rings has been generated [7]. In this material, the nonhexagonal rings alter the electronic properties of the nanoribbons due to the topological change. Similar to graphene nanoribbons, 2D nonbenzenoid carbon allotropes have also attracted great interest from researchers. In 2021, an article reported a 2D biphenylene network composed of $sp^{2}$-hybridized carbon atoms with periodically arranged four-, six-, and eight-membered rings [8], achieving a two-dimensional square-hexagon-octagon (SHO) lattice, as illustrated in Fig. 1. The SHO lattice as a new kind of unfrustrated 2D lattice offers a new platform for studying electronic correlations, magnetism, and superconducting properties.
@@
Figure 1: (a) The sketch of the SHO structure, constituted with hexagon unit cells. There are three types of bonds with different interactions from our definition: $J_{1}$ is indicated with black color, $J_{2}$ is red, and $J_{3}$ is blue. We also define the ratios between them: $\lambda_{1}=J_{2}/J_{1}$ and $\lambda_{2}=J_{3}/J_{1}$. A unit cell of the SHO lattice consists of six sites, which is shown in the black dashed square. (b) The lattice in square form is topologically equivalent to the SHO lattice.
@@
The electronic band structures were systematically investigated using density functional theory [9, 10, 11], and the possible superconductivity properties of biphenylene were studied using the random phase approximation [10]. Furthermore, the ground-state phase diagram and critical behavior of the Hubbard model on the SHO lattice were examined using the determinant quantum Monte Carlo method [12]. Additionally, the topological and electronic properties on an anisotropic SHO lattice were also studied [13]. However, the magnetic properties of the Heisenberg model on the SHO lattice remain an open question. It is of great interest to investigate the ground-state properties of the $S=1/2$ Heisenberg model on the SHO lattice. Although no SHO lattice magnetic materials have been identified in transition metals thus far, there is potential for their discovery in the future. Our research offers crucial theoretical analysis in anticipation of such findings. Moreover, magnetic systems of this type could potentially be realized in experimental platforms such as cold atoms or metal-organic frameworks. As a new theoretical model, it is of significant value for the systematic investigation of its rich phase diagram and critical behavior.
@@
In this work, we employ the stochastic series expansion (SSE) quantum Monte Carlo (QMC) [14, 15, 16, 17] and density matrix renormalization group (DMRG) [18, 19, 20, 21, 22, 23] methods to study the ground-state phase diagram of the $S=1/2$ Heisenberg model on the SHO lattice. To enrich the phase diagram and enhance its topological equivalence to the square lattice, we introduce the selected third-nearest-neighbor interaction $J_{3}$ along the $x$ direction. Notably, among the discovered phases—which include Néel, dimerized, and hexagon phases—we identify a Haldane-like phase [24, 25]. This phase has a unique ground state which does not break any symmetries and exhibits a finite triplet excitation gap. Its topological properties can be evidenced by the degeneracy of the ground-state energy levels under open boundary conditions (OBCs) and the twofold degeneracy of the entanglement spectrum. This phase is distinct and cannot be continuously connected to trivial phase in the presence of certain symmetries, making it a symmetry-protected topological (SPT) phase [26, 27, 28, 29]. Similar phases have been observed in $S=1/2$ two-leg ladder systems [30, 31, 32]. Furthermore, to establish connections with potential experimental material synthesis, we also investigate the phase diagram of this SPT phase with respect to other possible interactions. To obtain accurate phase boundaries and critical behaviors between different phases, we employ comprehensive finite-size scaling techniques [33, 34] to verify the universality class.
@@
The paper is organized as follows. In Sec. II, we introduce the model Hamiltonian and outline the physical quantities used to characterize the ground-state phase diagram of the SHO lattice. Then, in Sec. III, we employ SSE-QMC and DMRG methods combined with finite-size scaling to numerically determine the accurate phase diagram. Specifically, Sec. III.1 examines the case with $\lambda_{2}=J_{3}/J_{1}=0$, Sec. III.2 considers the scenario where $\lambda_{1}=J_{2}/J_{1}=0$, and Sec. III.3 explores the full phase diagram. Following that, Sec. III.4 focuses on the critical behavior of phase transitions on the SHO lattice. Moreover, Sec. III.5 investigates the stability of the SPT phase against other interactions. Finally, in Sec. IV, we summarize all identified ground-state phases and provide a concise discussion.
@@ II MODEL AND METHODS
The $S=1/2$ Heisenberg model on the SHO lattice is described by the Hamiltonian:
@@
$+J_{3}\sum_{\langle\langle\langle i,j\rangle\rangle\rangle}\mathbf{S}_{i}\cdot\mathbf{S}_{j}.$ (1)
Note that the SHO lattice and its topologically equivalent squarelike lattice are depicted in Figs. 1(a) and 1(b), respectively. The Hamiltonian includes several types of spin-spin interactions: $\langle i,j\rangle$ and $\langle i,j\rangle^{\prime}$ represent intra- and interhexagon nearest-neighbor couplings, respectively, while $\langle\langle\langle i,j\rangle\rangle\rangle$ corresponds to the selected third-nearest-neighbor interaction along the $x$ direction. This $J_{3}$ term also acts as a nearest-neighbor interaction in the topologically equivalent squarelike lattice. Here, $S_{i}$ denotes the spin operator at site $i$. For simplicity, we assume isotropic interhexagon interactions, meaning that the horizontal and vertical $J_{2}$ bonds are set equal. The $J_{3}$ coupling was introduced primarily to establish a structural connection with the conventional square-lattice Heisenberg model. To facilitate comparison of interaction strengths, we define the dimensionless ratios $\lambda_{1}=J_{2}/J_{1}$ and $\lambda_{2}=J_{3}/J_{1}$.
In the following text, unless specified, we generally employ periodic boundary conditions (PBCs) in QMC calculations. The linear system size is denoted as $L$, and the total number of grid points is given by $N=6L^{2}$. In the SSE-QMC calculation, we use the staggered magnetization along the $S^{z}$ direction and the squared staggered magnetization $M_{s}^{2}$ to illustrate the Néel order:
@@
$M_{s}^{2}=$ $\frac{1}{N^{2}}\sum_{ij}\phi_{i}\phi_{j}\langle\mathbf{S}_{i}\cdot\mathbf{S}_{j}\rangle.$ (3)
@@
$=\frac{3}{2\beta N}\left\langle\left(N_{a}^{+}-N_{a}^{-}\right)^{2}\right\rangle,\quad a=x,y,$ (5)
@@
Spin stiffness denotes the ground-state energy cost of applying a twisting angle $\theta$ to the rotors of a spin system, which can be calculated by the number of $S_{i}^{+}S_{j}^{-}$ and $S_{i}^{-}S_{j}^{+}$ operators in the SSE program. Here $E_{0}$ is the ground-state energy of the twisted Hamiltonian, $N$ is the number of spins in the system, $N_{a}^{+}$ is the number of $S_{i}^{+}S_{j}^{-}$ operators, and $N_{a}^{-}$ is the number of $S_{i}^{-}S_{j}^{+}$ operators. The value of $\beta=1/T$ is taken to be proportional to the system size $L$ when doing the scaling. The factor $3/2$ in the Binder cumulant is for normalization, i.e., $U_{2}\rightarrow 1$ when the system is a magnetic ordered state and $U_{2}\rightarrow 0$ when it is a disordered state, for $N\rightarrow\infty$.
When $J_{2}=0$, the system becomes a ladder system. To analyze the ground-state phases in this scenario, we use the DMRG method to compute the energy gaps and entanglement spectra. In our DMRG calculations, we use 4000 $SU(2)$ symmetric states to ensure that the density matrix truncation error is below $10^{-8}$. The energy gaps are defined as follows: The singlet gap $\Delta_{S}$ is the energy difference between the first excited state and the ground state in the $S=0$ sector. The triplet gap $\Delta_{T}$ is the energy difference between the lowest $S=1$ state and the $S=0$ ground state. The quintuplet gap $\Delta_{Q}$ is the energy difference between the lowest $S=2$ state and the $S=0$ ground state, i.e.,
@@ III NUMERICAL RESULTS
Using SSE-QMC method and finite-size scaling, we obtain the accurate ground-state phase diagram of the $S=1/2$ Heisenberg model on the 2D SHO lattice, which is shown in Fig. 2. There are four gapped phases, namely the hexagon phase, the orthogonal staggered dimer (OSD) phase, the ladder staggered dimer (LSD) phase, and the Haldane-like SPT phase, surrounding the gapless Néel antiferromagnetic (AFM) phase. In the subsequent subsection, we will elaborate on the method used to obtain the phase diagram and provide a detailed overview of the distinct phases.
@@
Figure 2: The ground-state phase diagram of the $S=1/2$ Heisenberg model on the 2D SHO lattice. In total there are five phases in the phase diagram, namely the hexagon phase, the orthogonal staggered dimer (OSD) phase, the ladder staggered dimer (LSD) phase, the Haldane-like SPT phase, and the Néel antiferromagnetic (AFM) phase.
III.1 $J_{1}\text{-}J_{2}$ ground-state phase diagram
@@
First, we investigate the ground state of the system with $\lambda_{2}=J_{3}/J_{1}=0$ fixed and a changing $\lambda_{1}=J_{2}/J_{1}$. Initially, we examined the ground state of the Heisenberg model on the SHO lattice at two extremes: $\lambda_{1}$ approaching zero and positive infinity. At $\lambda_{1}=\lambda_{2}=0$, the model reduces to isolated, orderly hexagonal lattices. The ground state here is a direct product of hexagonal singlet states with finite local excitation gaps. After adding weak interhexagon interactions, the system is in a hexagon phase with gapped lowest triplet excitation. When $\lambda_{1}\to\infty$ and $\lambda_{2}=0$, the ground state adiabatically connects to a vertically and horizontally regular dimer product state, termed the orthogonal staggered dimer phase.
@@
Regarding the phase between the hexagon phase and the OSD phase, given the even-number sublattice structure of the SHO lattice, it is possible that a long-range magnetic order exists. And it is tentatively speculated to be the Néel AFM phase. By extrapolating the squared magnetization at $\lambda_{1}=1$ [as shown in Fig. 3 (b)], we can confirm that the intermediate phase is indeed a Néel AFM phase. To characterize the phase boundaries among the three phases, we show the intersection of Binder cumulant $U_{2}$ for different sizes when $\lambda_{1}\in[0,3]$, which is illustrated in Fig. 3 (a). To pinpoint the crossing location precisely, we computed additional, denser data points in the vicinity of the intersection, enabling a precise determination of the crossing point, as shown in the inset of Fig. 3 (a). Subsequently we perform an extrapolation of this intersection to the thermodynamic limit of $1/L\to 0$ using scaling function $\lambda_{1,c}(L)=\lambda_{1,c}(\infty)+aL^{-\omega}$ [14]. Based on the finite-size scaling results, it is evident that the system undergoes phase transitions at $\lambda_{1,c}=0.664(1)$ and $1.379(1)$ in the thermodynamic limit.
@@
Figure 3: (a) Variation of Binder cumulant $U_{2}$ with $\lambda_{1}$ at different sizes when $\lambda_{2}=0$. The inset in (a) is the extrapolation of the two intersection positions of the Binder cumulant as a function of the size, using the extrapolation function in the form of $\lambda_{1,c}(L)=\lambda_{1,c}(\infty)+aL^{-\omega}$ [14]. (b) Second-order polynomial extrapolation of $M_{s}^{2}$ as a function of $L$ when $\lambda_{1}=1$. The nonzero extrapolated value indicates that there is a nonzero staggered magnetic order in the thermodynamic limit: $M^{2}_{s}(L\to\infty)=0.0609(2)$, and $SU(2)$ continuous symmetry breaking occurs at the intermediate phase.
@@
Hence, when $\lambda_{2}=0$ and $\lambda_{1}$ varies (i.e., in the $J_{1}\text{-}J_{2}$ model), the ground-state phase diagram consists of three phases: the hexagon phase, the AFM phase, and the OSD phase. The hexagon phase and OSD phase are nonmagnetic, featuring nonzero triplet excitation energy gaps. In contrast, the AFM phase is magnetic, characterized by gapless Goldstone mode excitations.
III.2 $J_{1}\text{-}J_{3}$ ground-state phase diagram